OpenStudy (anonymous):

What's the difference between f(a^2) and [f(a)]^2?

OpenStudy (anonymous):

an example will help you :) ! f(x)=2x+1 what's f(a^2).....and [f(a)]^2

OpenStudy (anonymous):



OpenStudy (anonymous):

yes that was f(a^2) now [f(a)]^2 ?

OpenStudy (anonymous):

2x+1?

OpenStudy (anonymous):

f(x)=2x+1 f(a)=2a+1 [f(a)]^2=.........

OpenStudy (australopithecus):

\[[f(a)]^2 = (2a + 1)^2\] \[f(x^2) = 2x^2 + 1\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Yw ! did you the difference ?

OpenStudy (kinggeorge):

Just a quick interjection here, but are you sure that \[[f(x)]^2=f(x)\cdot f(x)\]And not\[[f(x)]^2=f(x)\circ f(x)=f(f(x))?\]I think this is rather ambiguous. If you aren't sure however, I'd stick with the first way, which Australopithecus described above.